### v1

# A Statistical Test for Comparing the Linkage and Admixture Model Based on Central Limit Theorems

Abstract
In the Admixture Model, the probability that an individual carries a certain allele at a specific marker depends on the allele frequencies in $K$ ancestral populations and the proportion of the individual’s genome originating from these populations. The markers are assumed to be independent. The Linkage Model is a Hidden Markov Model (HMM) that extends the Admixture Model by incorporating linkage between neighboring loci.
This study investigates the consistency and central limit behavior of maximum likelihood estimators (MLEs) for individual ancestry in the Linkage Model, complementing earlier results by (Pfaff et al., 2004; Pfaffelhuber and Rohde, 2022; Heinzel, 2025) for the Admixture Model. These theoretical results are used to prove theoretical properties of a statistical test that allows for model selection between the Admixture Model and the Linkage Model. Finally, we demonstrate the practical relevance of our results by applying the test to real-world data from The 1000 Genomes Project Consortium (2015).
Carola Sophia Heinzel
Department of Mathematical Stochastics,
Ernst-Zermelo Straße 1, Freiburg im Breisgau, 79140,Germany
carola.heinzel@stochastik.uni-freiburg.de
Keywords: Linkage Model, Admixture Model, Central Limit Results, Maximum Likelihood Estimator, Consistency, Statistical Test, Model Selection
1 Introduction
Both the Linkage Model (Falush et al., 2003) and the Admixture Model (e.g. Alexander et al. (2009)) are widely used to explain the genetic data of individuals. They are similar in that both assume that genetic data can be described by the ancestry proportions $q$ of an individual from $K$ ancestral populations and the allele frequencies in these populations. However, the Linkage Model additionally incorporates the distance between loci through a Hidden Markov Model (HMM). Furthermore, in the Linkage Model, there exists a parameter $r$, which can be interpreted as the number of generations since an admixture event (Falush et al., 2003). As a result, Falush et al. (2003) formulated the Linkage Model as a Hidden Markov Model (HMM). The Admixture Model can be seen as a special case of the Linkage Model with $r=\infty$, i.e. the data is assumed to be independent across markers.
A natural question that arises is which model fits a given dataset better. This leads to a nested model selection problem as described by Anderson and Burnham (2004). In our case, we consider the statistical hypothesis test:
$H_{0}:r=\infty\quad\text{vs.}\quad H_{1}:r\in[0,\infty).$ (1)
This is a classical nested test problem, for which asymptotic theory exists (Wilks, 1938). However, the test is only valid as an asymptotic level-$\alpha$ test if the MLEs for both ancestry and $r$ are asymptotically normally distributed under the Linkage Model. Establishing asymptotic normality requires results from the theory of HMMs.
Consistency of MLEs in HMMs with finite state and observation spaces has been addressed in foundational work by Baum and Petrie (1966) and Petrie (1969). More recent work has relaxed several of the original assumptions, e.g., (Leroux, 1992; Douc et al., 2004, 2011; Genon-Catalot and Laredo, 2006; Le Gland and Mevel, 2000a, b), but these results assume time-homogeneous Markov chains. Central limit theorems for MLEs in HMMs have also been developed (Bickel et al., 1998; Douc et al., 2004; Jensen and Petersen, 1999; Brouste and Kleptsyna, 2010). Overviews of the statistical theory of HMMs can be found in Ephraim and Merhav (2002); Cappé et al. (2005). However, the specific case of a time-inhomogeneous Markov chain that maintains the same stationary distribution across all time points—as occurs in the Linkage Model when $q$ is the initial distribution—has not yet been studied.
Several statistical tests and heuristics have been proposed for model selection in genetics. For example, to decide whether a pair of loci is suitable for STRUCTURE (Pritchard et al., 2000), Kaeuffer et al. (2007) suggested using $r_{LD}$, a measure of linkage disequilibrium introduced by Hill and Robertson (1968). They ran STRUCTURE with the Linkage Model and used logistic regression to assess the impact of $r_{LD}$ on the detection of population structure. Other studies have examined the sensitivity of cluster identification to genetic and ecological constraints (Rosenberg et al., 2001; Evanno et al., 2005).
To achieve our goal of establishing the theoretical properties of the test in (1), we proceed as follows: First, we precisely define both the Admixture Model and the Linkage Model. Next, we prove the asymptotic normality of the MLEs in the Linkage Model, both when the number of chromosomes and when the number of markers tends to infinity. Based on these results, we construct a statistical test for (1) and prove that it is an asymptotic level-$\alpha$ test. We also provide a method to quantify the uncertainty of the MLEs in the Linkage Model. Finally, we evaluate the statistical test through simulations and apply it to real data from The 1000 Genomes Project Consortium (2015).
2 Models
We first define the Linkage Model for one individual and bi-allelic markers.
Definition 2.1 (Linkage Model for Haploid Individuals).
The number of alleles at chromosome $c\in\{1,...,C\}$ at marker $m\in\{1,...,M_{c}\}$ is called $X_{c,m}\in\{0,1\}.$ We denote the genetic distance in centi Morgan (cM) between the loci m1m-1 and $m$ on chromosome $c$ by $d_{c,m}$. Let $\mathbb{S}^{K}$ be the (K1)(K-1)-dimensional simplex and let $q:=(q_{1},\ldots,q_{K})\in\mathbb{S}^{K}$ be the ancestries of the individual from population 1,,K.1,...,K. The frequency of an allele in population $k\in\{1,...,K\}$ at marker $m$, chromosome $c$ is called $p_{c,k,m}.$ The random variable $Z_{c,m},m=1,...,M_{c},c=1,...,C$ names the ancestral population of the allele at marker $m$ at chromosome $c$. Based on this, we define the Markov process
$\mathbb{P}_{q,r}(Z_{c,1}=k)$ $=q_{k},$
$\mathbb{P}_{q,r}(Z_{c,m}=\tilde{k}|Z_{c,m-1}=k)$ $=\begin{cases}e^{-d_{c,m}r}+\left(1-e^{-d_{c,m}r}\right)q_{k},\textup{ if }k=\tilde{k}\\ \left(1-e^{-d_{c,m}r}\right)q_{k},\textup{ else},\end{cases}$ (2)
$\mathbb{P}_{q,r}(Z_{c+1,1}=k|Z_{c,M_{c}})$ $=q_{k}.$
The emission probability is defined by
$\mathbb{P}_{q,r}(X_{c,m}=x|Z_{c,m})=\mathbb{P}_{q,r}\left(\textup{Ber}\left(q_{Z_{c,m}}p_{c,Z_{c,m},m}\right)=x\right).$
We define the log-likelihood, for $M_{total}:=\sum_{c=1}^{M_{C}},$
$\ell((x_{1,1},...,x_{C,M}),(q,r)):=\frac{\log\left(\mathbb{P}_{q,r}((X_{c,m}=x_{c,m})_{c=1,...,C,m=1,...,M_{c}}))\right)}{M_{total}}.$
We write $(q^{0},r^{0})$ for the true parameters $(q,r)$ and $\mathbb{E},\mathbb{P}$ for the expected value and the probability respectively, with respect to the true parameters $q^{0},r^{0}.$
We always assume that $d_{c,m}\geq 0$ for all $m\in\{1,\ldots,M\}$ and that the allele frequencies and $d_{c,m}$ are known.
Remark 2.2 (Diploid Case).
Let $Z^{j}:=(Z^{j}_{c,1},...,Z^{j}_{c,M}),j=1,2,$ be two independent Markov chains with transition matrix (2). Let $X_{c,m}^{dip}$ be the number of alleles at chromosome $c,$ marker $m$ and let $x\in\{0,1,2\}.$ The emission probabilities for the diploid case are defined by
$\mathbb{P}_{q,r}\left(X^{dip}_{c,m}=x|Z^{1}_{c,m},Z^{2}_{c,m}\right)$
$\hskip 28.45274pt=\begin{cases}q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m}\cdot q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m},&x=2\\ q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m}(1-q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m})+q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m}(1-q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m}),&x=1\\ (1-q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m})(1-q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m}),&x=0.\end{cases}$
For the diploid case, the information about the maternal and the paternal copies (Choi et al., 2018), i.e. phased data, is important. There are two kinds of approaches to receive phased data: laboratory-based methods (Zheng et al., 2016; Amini et al., 2014; Duitama et al., 2012) and computational methods (Choi et al., 2018). Examples for the latter type are (Snyder et al., 2015; Loh et al., 2016; Delaneau et al., 2012). Falush et al. (2003) also proposed a method to deal with linkage without phasing. However, in this study, we assume that we have phased data as e.g. in The 1000 Genomes Project Consortium (2015).
While phased data is important for the Linkage Model, the Admixture Model can deal with unphased data without any problems.
Definition 2.3 (Admixture Model).
The Admixture Model is a special case of the Linkage Model with $r=\infty.$
Extending the models to a general number of individuals and markers with arbitrary number of alleles is straightforward.
We do not only need to know about the Linkage Model, but we also need some general notation is mentioned in Notation 2.4.
Notation 2.4 (General Notation).
For a positive definite matrix $A$, we write $A\succ 0.$ Additionally, we write $X_{i:i+s}$ for $X_{i},...,X_{i+s}.$
3 Theoretical Main Results
In this chapter, we prove the main results, i.e. consistency and central limit results, if the number of markers, $M_{total},$ tends to infinity. Therefore, we need assumption 3.1.
Assumption 3.1.
We assume that
  • (A1)
    the parameter space $\Theta:=\{[\kappa_{q},\kappa_{q}^{\prime}]^{K-1}\times[r_{lb},\infty]\}\subset\mathbb{R}^{K+1}$ is a compact set.
  • (A2)
    it holds $1>\kappa_{p}^{\prime}\geq p_{\cdot,m}\geq\kappa_{p}>0$ for all $m\in\{1,...,M\}.$
  • (A3)
    $p_{k,m}\neq p_{\ell,m}$ for infinitely many markers m and for all $k,\ell\in\{1,..,K\},k\neq\ell.$
  • (A4)
    Let $\lambda_{c,m}$ be the second largest of the transition matrix $(\mathbb{P}(Z_{c,m}=k|Z_{c,m-1}=k^{\prime}))_{k=1,...,K,k^{\prime}=1,...,K}$. It holds
    $\prod_{c=1}^{C}\prod_{m=1}^{M_{c}}\lambda_{c,m}\xrightarrow[]{M_{total}\to\infty}0$
  • (A5)
    There exists a constant $\kappa_{d}>0$, such that $d_{c,m}\geq\kappa_{d}$ for all $m\in\{1,...,M_{c}\},c=1,...,C.$
Figure 1 gives an overview of the results in this paper.
ConsistencyCLTAsymptotic Distribution of the Test Statistic
Figure 1: Overview of results in this section and how they are connected to each other.
We also state the theory under the following conditions concerning the data.
Remark 3.2 (Assumptions).
We assume that the data is haploid and that we only have bi-allelic markers. Additionally, we only consider one individual. However, extending the theory to more general cases is straightforward.
We start with the consistency of the MLE (Theorem 1).
Theorem 1 (Consistency of the MLE).
Let Assumption 3.1 hold. Then, for the MLE
$\left(\hat{Q}^{C,M},\hat{R}^{C,M}\right):=\textup{argmax}\{(q,r)\mapsto\ell((x_{1,1},...,x_{C,M}),(q,r))\}$
it holds
$\mathbb{P}\left(\lim_{M_{total}\to\infty}\bigg{|}\left(\hat{Q}^{C,M},\hat{R}^{C,M}\right)-\left(q^{0},r^{0}\right)\bigg{|}\geq\epsilon\right)=0$
for any $\epsilon>0.$
Based on Theorem 1, we can infer a CLT for the MLE, if the true parameters $\left(q^{0},r^{0}\right)$ are in the interior of the parameter space.
Theorem 2 (Central Limit Theorem for the MLE).
We define the Fischer information
$J_{q^{0},r^{0}}:=\lim_{M_{total}\to\infty}\mathbb{E}\left(\frac{\partial^{2}}{\partial(q,r)^{2}}\ell\left((x_{1,1},...,x_{C,M}),(q,r))\big{|}_{(q,r)=\left(q^{0},r^{0}\right)}\right)\right)$
and assume $J_{q^{0},r^{0}}\succ 0$. Let the true parameter be in the interior of the parameter space. Then, it holds
$\sqrt{M_{total}}\left((\hat{Q}^{C,M},\hat{R}^{C,M})-(q^{0},r^{0})\right)\xRightarrow{M_{total}\to\infty}\mathcal{N}\left(0,J_{q^{0},r^{0}}^{-1}\right).$
Remark 3.3.
Douc (2005) has proven the invertibility of the Fischer Information under certain constraints for stationary HMMs, i.e. in our case, if the allele frequencies and the distance between the markers are identical for every marker. This also means that the markers are all at different chromosomes or at the same chromosome.
According to Wilks (1938), since the MLE in the Linkage Model is asymptotically normally distributed, Theorem 3 holds.
Definition 3.4 (Statistical Test).
We define the test statistic for the test (1) by
$\Lambda:=-2\ln\left(\frac{\textup{argmax}\{(q,r)\mapsto\ell((X_{1,1},...,X_{C,M}),(q,\infty))\}}{\textup{argmax}\{(q,r)\mapsto\ell((X_{1,1},...,X_{C,M}),(q,r)):(q,r)\in\Theta\}}\right).$
Furthermore, let $\chi^{2}_{1-\alpha}$ be the $1-\alpha$-quantile of the $\chi^{2}(1)$-distribution. We reject $H_{0}$, if $\Lambda>\chi^{2}_{1-\alpha}.$
Theorem 3 (Asymptotic distribution of the test statistic).
It holds
$\Lambda\xRightarrow{M_{total}\to\infty}\chi^{2}(1)$
under the null hypothesis of (1).
4 Application to Data
In this chapter, we first evaluate the performance of the test from Definition 3.4 by using simulated data. Afterwards, we apply the test to data from The 1000 Genomes Project Consortium (2015) and compare the uncertainty of the MLEs in the Linkage Model to the ones in the Admixture Model.
4.1 Simulated Data
To evaluate the statistical test by using the type 1 and type 2 error, we simulate the data according to the Linkage Model and the Admixture Model with different values for $M,r$ and $(d_{c,m})_{c=1,...,C,m=1,...,M_{c}-1}.$ More precisely, we simulated 100 markers on one single chromosome with $d_{1,m}=d_{1,m+1}\forall m,d:=d_{1,m}\in\{0.1,0.5,1,2,5,10\}$ and $r\in\{1,10,100\}.$ We choose $0.05$ as the significance level. We calculated the type 1 and the type 2 error by repeating the experiment 100 times per possible combination of $r$ and $d$.
The results are shown in Figure 2. Especially the type 1 error is small and lower than 0.05. Furthermore, for increasing $r$ and $d$ the type 1 and the type 2 error tend to increase. This is not surprising since a large $r$ corresponds to a high similarity between the two models. To summarize, the simulation results show that the test has both, a high power and a small type 1 error.
Figure 2: Evaluation of the statistical test for simulated data.
4.2 Real Data
We also applied our theoretical results to the data from The 1000 Genomes Project Consortium (2015) with the AIM set by Kidd et al. (2014) that consists of $M_{total}=55$ markers on 20 different autosomes ($C=20$). We always used all 2504 individual except the one that we are testing to calculate the allele frequencies as relative frequency. The values for $d$ for the AIM set by Kidd et al. (2014) in cM are shown in Figure 3.
Figure 3: Genetic Distances of the markers in the AIM set by Kidd et al. (2014). We rounded them to integer values.
The distance between the loci at a single chromosome was between below 1 cM and approximately 129 cM. We used individuals from Africa (AFR), Europe (EUR), South-East-Asian (SAS), East-Asian (EAS) and Admixed Americans (AMR), i.e. $K=5$.
The results of the statistical test are shown in Figure 4. Altogether, for 312 out of 2504 ($\approx 12.5\%$ of the) individuals, the null hypothesis that the Admixture Model fits better to the data, cannot be rejected. Thereby, we see strong differences between the populations: For people from Africa and East Asian, the fraction of individuals for which the null hypothesis is rejected is much smaller than for the other populations. This emphasizes that it is not sufficient to name a general rule to decide for either the Linkage Model or the Admixture Model that only depends on the marker set. Instead, a method – like the statistical test (1) – is required, that decides for every individual or population separately.
Figure 4: Results of the statistical test.
We also compare the covariance for both, the MLE in the Linkage Model and the MLE in the Admixture Model to each other for individual HG00096 from Great Britain. Therefore, we again used the AIM Set by Kidd et al. (2014). The results are presented in figures 6 and 6. There, we see that the variance of the estimators in the Linkage Model are higher than in the Admixture Model. This aligns with our expectations since the Linkage Model contains more parameters. Although both models clearly identify EUR as the primary ancestral population, the estimated values differ significantly.
Figure 5: Covariance Matrix for the MLE in the Admixture Model. We considered individual HG00096. The MLE for $q$ was $(1,0,0,0,0).$
Figure 6: Covariance Matrix of the MLE in the Linkage Model. We considered individual HG00096. The MLE for $q$ was $(0.8,0.04,0.04,0.08,0.04)$ and for $r$ approximately $1.8.$
Remark 4.1 (Test for a whole population).
We also provide the code for testing whether the Linkage Model or the Admixture Model fits better to the data of $N$ individuals on GitHub. Therefore, we could either assume that $r$ depends on the individual, i.e. we would have $NK$ parameters that we have to estimate or that we have only one global parameter $r,$ i.e. we estimate N(K1)+1N(K-1)+1 parameters. We decided for the latter case. However, the allele frequencies are supposed to be known in this case. Additionally, the diploid case is also implemented.
5 Proofs of the Main Results
We first prove that the MLE is asymptotically unique. Based on this, we prove consistency and central limit results for the MLE in the Linkage Model. For all proofs, we assume $C=1$ to simplify the notation.
To prove the consistency and the CLT, we represent the log-likelihood as a sum, i.e. it holds
$\ell((X_{1},...,X_{M}),(q,r))$
$\quad=\frac{1}{M}\sum_{m=1}^{M}\underbrace{\log\left(\int\mathbb{P}_{q,r}(X_{m}|Z_{m}=z)\mathbb{P}_{q,r}(Z_{m}=dz|X_{1},...,X_{m-1})\right)}_{=:D_{m}^{q,r}}$
according to van Handel (2008), Proposition 6.4, for the homogeneous case. This can easily extended to the inhomogeneous case.
5.1 Unique MLE
In this section, we use that the limit
$\ell((X_{1},...,X_{M}),(q,r))\xrightarrow[]{M\to\infty}\ell(q,r)$
exists almost surely. We will prove this in section 5.2.
Proving the uniqueness of a MLE for HMM has been already considered extensively for a finite state space, see e.g. (Finesso, 1990; Gilbert, 1959; Blackwell and Koopmans, 1957; Petrie, 1969). In this section, we aim to prove the asymptotic uniqueness of the MLE in the Linkage Model, i.e. we aim to prove Theorem 4.
Theorem 4.
Let Assumption 3.1 hold. Then, it holds
$\ell(q^{0},r^{0})>\ell(q,r)\Leftrightarrow q,r\neq q^{0},r^{0}.$ (3)
Therefore, we first prove that the parameters are identifiable based on a finite number of observations. Based on this, we prove that if $\ell(q^{0},r^{0})\neq\ell(q,r)$, then the measures $\mathbb{P}_{q,r}|_{(X_{m})_{m=1,2,...}}$ and $\mathbb{P}_{q^{0},r^{0}}|_{(X_{m})_{m=1,2,...}}$ are mutually different. Based on this, we can then conclude the claim.
We write $(q,r)\sim(q^{\prime},r^{\prime})$, if there exists an invariant distribution $\pi,\pi^{\prime}$ such that $\mathbb{P}_{(q,r)}^{X}=\mathbb{P}_{(q^{\prime},r^{\prime})}^{X}$ holds. Therefore, we only consider the case $K=2$ as the proof involves naming conditions that yield to the full rank of the emission matrix.
Lemma 5.1 (Identifiability of the Parameters).
Let Assumption 3.1 hold and let $K=2.$ If either the allele frequencies $p_{\cdot,i},...,p_{\cdot,i+s}$ are linearly independent and $s\geq 4$ or $s=3$ and $q\neq p_{1,2}/(p_{1,2}+p_{2,2})$ holds and in addition $p_{1,m}\neq p_{2,m}\forall m\in\{i,...,i+s\}$, the parameters $(q,r)$ are identifiable based on the random variables $X_{i},...,X_{i+s}.$
Proof.
To save some notation, we write $A_{i}(z_{i},z_{i+1})=\mathbb{P}(Z_{i+1}=z_{i+1}|Z_{i}=z_{i}),B_{j}(x_{j},z_{j})=\mathbb{P}(X_{j}=x_{j}|Z_{j}=z_{j}).$ In order to apply Kruskal’s Theorem, we first need a decomposition of the marginal density of $X_{1},X_{2},X_{3}.$ Let $M_{j}$ be a stochastic matrix, where the $i^{th}$ row is denoted by $m^{j}_{i}.$ We define
$[M_{1},M_{2},M_{3}]_{x_{1},x_{2},x_{3}}=\sum_{i=1}^{2}m^{1}_{i}(x_{1})m^{2}_{i}(x_{2})m^{3}_{i}(x_{3})$
and $M_{1},M_{2},M_{3}]$ is a three-dimensional array for which the element $(x_{1},x_{2},x_{3})$ equals $M_{1},M_{2},M_{3}]_{x_{1},x_{2},x_{3}}.$ It holds
$\mathbb{P}(X_{1},X_{2},X_{3})$ $=\sum_{z_{1},z_{2},z_{3}}\pi_{z_{1}}A_{1}(z_{1},z_{2})A_{2}(z_{2},z_{3})B_{2}(X_{2},z_{2})B_{3}(X_{3},z_{3})B_{1}(X_{1},z_{1})$
$=[\underbrace{\pi B_{1}A_{1}}_{=:M_{1}},B_{2},\underbrace{A_{2}B_{3}}_{=:M_{2}}].$
The proof is divided into two steps:
  • (i)
    prove that the entries of $M_{1},M_{2},B_{2}$ are identifiable from the marginal density of $X_{1},X_{2},X_{3}$. Therefore, we use theory by Allman et al. (2009); Kruskal (1977).
  • (ii)
    conclude that the parameters $q,r$ are identifiable.
We start with (i). Therefore, we have to prove $rk(M_{1})=rk(B_{2})=rk(M_{2})=2.$ It holds
$\det(B_{2})=\bigg{|}\begin{pmatrix}q_{1}p_{1,2}&(1-q_{1})p_{2,2}\\ 1-q_{1}p_{1,2}&1-(1-q_{1})p_{2,2}\end{pmatrix}\bigg{|}=q_{1}p_{1,2}-(1-q_{1})p_{2,2}.$
Hence, $B_{2}$ has the full rang if and only if $q_{1}\neq\frac{p_{1,2}}{p_{1,2}+p_{2,2}}.$ This is the reason, why we need the condition about the linear independence of the allele frequencies (since we can apply Kruskal’s theorem to two different time points within the interval). We have
$M_{1}=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$
with entries
$a$ $=(1-q_{1}p_{1,1})\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right)+\left(1-(1-q_{1})p_{2,1}\right)(1-e^{-rd_{1}})(1-q_{1})$
$=(1-q_{1}p_{1,1})(q_{1}+e^{-rd_{1}}(1-q_{1}))+(1-p_{2,1}+q_{1}p_{2,1})(1-q_{1})(1-e^{-rd_{1}})$
$b$ $=q_{1}p_{1,1}\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right)+p_{2,1}(1-q_{1})^{2}(1-e^{-rd_{1}}),$
$c$ $=(1-q_{1}p_{1,1})(1-e^{-rd_{1}})(1-q_{1})+\left(1-(1-q_{1})p_{2,1}\right)\left((1-e^{-rd_{1}})(1-q_{1})+e^{-rd_{1}}\right),$
$d$ $=q_{1}p_{1,1}(1-e^{-rd_{1}})(1-q_{1})+p_{2,1}(1-q_{1})\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right).$
It holds
$\det(M_{1})=e^{-rd_{1}}(1-e^{-rd_{1}})q_{1}(1-q_{1})(p_{1,1}+p_{2,1}-2p_{1,1}p_{2,1})=0$
$\Leftrightarrow(r=\infty)\lor(r=0)\lor(q_{1}\in\{0,1\})\lor(p_{1,1}=p_{2,1}=0)\lor(p_{2,1}=p_{1,1}=1).$
Here, we excluded every case except the case $r=\infty.$ This case leads to the Admixture Model which has already been discussed by Heinzel (2025). We additionally calculate
$\mathbb{P}(X_{k-2}=x\mid Z_{k-1}=z)$ $=\sum_{z^{\prime}}\mathbb{P}(X_{k-2}=x,Z_{k-2}=z^{\prime}\mid Z_{k-1}=z)$
$\propto\sum_{z^{\prime}}\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=z^{\prime})\cdot\mathbb{P}(Z_{k-2}=z^{\prime}\mid Z_{k-1}=z)$
$\propto\sum_{z^{\prime}}\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=z^{\prime})\cdot\mathbb{P}(Z_{k-1}=z^{\prime}\mid Z_{k-2}=z)$
$\propto\underbrace{\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=0)}_{=(q_{1}p_{1,3})^{x}(1-q_{1}p_{1,3})^{1-x}}\cdot\mathbb{P}(Z_{k-1}=0\mid Z_{k-2}=z)$
$\hskip 28.45274pt+\underbrace{\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=1)}_{=(q_{2}p_{2,3})^{x}(1-q_{2}p_{2,3})^{1-x}}\cdot\underbrace{\mathbb{P}(Z_{k-1}=1\mid Z_{k-2}=z)}_{=1_{z=1}e^{-rd_{2}}+q_{z}(1-e^{-rd})}$
$\propto(q_{1}p_{1,3})^{x}(1-q_{1}p_{1,3})^{1-x}\left(1_{z=0}e^{-rd_{2}}+q_{z}(1-e^{-rd})\right)$
$\hskip 28.45274pt+(q_{2}p_{2,3})^{x}(1-q_{2}p_{2,3})^{1-x}\left(1_{z=1}e^{-rd_{2}}+q_{z}(1-e^{-rd})\right).$
This leads to
$M_{2}=\begin{pmatrix}a_{2}&b_{2}\\ c_{2}&d_{2}\end{pmatrix}$
with entries
$a_{2}$ $=(1-q_{1}p_{1,3})\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right)+(1-(1-q_{1})p_{2,3})q_{1}(1-e^{-rd_{2}})$
$b_{2}$ $=(q_{1}p_{1,3})\left(e^{-rd_{2}}+q_{1}(1-e^{-rd_{2}})\right)+((1-q_{1})p_{2,3})q_{1}(1-e^{-rd_{2}}),$
$c_{2}$ $=(1-q_{1}p_{1,3})(1-q_{1})(1-e^{-rd_{2}})+(1-(1-q_{1})p_{2,3})\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right),$
$d_{2}$ $=(q_{1}p_{1,3})(1-q_{1})(1-e^{-rd_{2}})+((1-q_{1})p_{2,3})\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right).$
Hence, it holds $rk(M_{2})=1\Leftrightarrow p_{2,3}=p_{1,3},q_{1}=0.5.$ Now, we come to step (ii), i.e. we have to ensure that the parameters are uniquely determined based on the entries. For the emission probabilities, this is a direct consequence of the assumption $p_{k,m}\neq p_{\ell,m}.$ Then, since $d_{m}>0,$ it is also immediate to infer the value $r$ from the other parameters.
Remark 5.2 (Identifiability in the Admixture Model for $K=2$).
Heinzel (2025) did not consider the identifiability for finitely many markers in the Admixture Model. For $K=2$, it holds
$\mathbb{P}(X_{m}=x)$ $=\langle q,p_{\cdot,m}\rangle^{x}(1-\langle q,p_{\cdot,m}\rangle)^{1-x}$
$=(q_{1}p_{1,m}+(1-q_{1})p_{2,m})^{x}(1-q_{1}p_{1,m}-(1-q_{1})p_{2,m})^{1-x}$
$=\begin{cases}p_{2,m}+q_{1}(p_{1,m}-p_{2,m})=:c_{m},x=1\\ 1-p_{2,m}+q_{1}(-p_{1,m}+p_{2,m}),x=0\end{cases}.$
This leads to
$q_{1}=\frac{c_{m}-p_{2,m}}{p_{1,m}-p_{2,m}},$
if $p_{1,m}=p_{2,m}.$ Hence, $q_{1}$ is uniquely determined if and only if $p_{1,m}=p_{2,m}.$
Remark 5.3.
In the latter, we will – without a proof – just assume that the parameters are identifiable from the marginal density of the observations. The proof is basically the same. However, we of course have to increase the number of observations form $3$ for $K=2$ to 2(K1)+2.2(K-1)+2. Additionally, we then have to verify again that the matrices of the decomposition of the marginal density have the full rank under certain conditions.
Now, we aim to infer (3) from the identifiability of the parameters. Therefore, we first prove a law of large numbers. Then, based on this, we conclude (3). This is done by using the following particularity of the hidden chain (Lemma 5.4) and that $(Z_{m})_{m=1,2,...}$ is uniformly ergodic (Lemma 5.6).
Lemma 5.4 (Invariant Measure).
For each $m\in\mathbb{N},$ the probability measure $q$ is the unique invariant measure for the transition matrix of $Z_{m}$.
Definition 5.5 (Uniformly Ergodic).
Let $|\cdot|_{TV}$ be the total variation norm. A Markov chain with transition matrices $T_{i}^{\theta}$ is called uniformly ergodic, if it holds
$|T^{\theta}_{n}...T^{\theta}_{1}-\pi|_{TV}\leq c_{\theta}\rho_{\theta}^{n}$
for $\rho_{\theta}<1,c_{\theta}\in\mathbb{R}.$
Lemma 5.6 (Uniformly Ergodicity for the Linkage Model).
Under Assumption 3.1, $(Z_{m})_{m=1,2,...}$ is uniformly ergodic.
Proof.
This claim is derived from Theorem 3.3 in Saloff-Coste and Zúniga (2007). ∎
Now, we are ready to name an LLN for the Linkage Model.
Lemma 5.7.
Let Assumption 3.1 hold and let $f_{i}$ be uniformly bounded by $f_{\infty}$. Then, it holds
$\mathbb{P}^{\nu}_{\theta}\left(\frac{1}{n}\left|\sum_{i=1,...,n}f_{i}(X_{i:i+s})-\pi(f_{i})\right|\geq\epsilon\right)\leq C\exp\left(-\frac{n}{8f_{\infty}}\right)$
for a constant $C>0$ and for any probability measure $\nu.$
Proof.
Without loss of generality, we assume $\mathbb{E}_{\pi}(f_{i}(X_{i:i+s}))=0$. From Lemma 5.6, we know that the Markov chain $(Z_{m})_{m=1,2,...}$ is uniformly ergodic. The proof is similar to Theorem 14 in Douc (2005): We first write
$\xi_{i,j}:=\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|Z_{0:i+j},X_{0:i+j})+\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|Z_{0:i+j-1},X_{0:i+j-1}).$
Then, since
$\sum_{j=0}^{s}\sum_{i=1}^{n}\xi_{i,j}$ $=\sum_{i=1}^{n}\sum_{j=0}^{s}\xi_{i,j}=\sum_{i=1}^{n}\xi_{i,s}-\xi_{i,0},$
it holds
$\sum_{i=1,...,n}f_{i}(X_{i:i+s})=\sum_{j=0}^{s}\sum_{i=1}^{n}\xi_{i,j}+\sum_{i=1}^{n}\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|Z_{0:i-1},X_{0:i-1}).$
Now, we can apply Azuma-Hoeffding inequality, which leads to
$\mathbb{P}^{\nu}\left(\bigg{|}\sum_{i=1}^{n}\xi_{i,j}\bigg{|}\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{8nf_{\infty}}\right).$
Now, we note that $\sum_{i=1}^{n}\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|X_{0:i-1},Z_{0:i-1})=F(Z_{i-1})$ for a certain function $F.$ We now prove that
$\mathbb{P}\left(\bigg{|}\sum_{i=1}^{n}\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|X_{0:i-1},Z_{0:i-1})\bigg{|}\geq t\right)2\leq\exp\left(-\frac{t^{2}}{8nf_{\infty}}\right).$
This is Theorem 17 in Douc et al. (2011).
To prove this claim, we apply Etemadi’s inequality and Bernstein’s inequality to the $\xi_{k}$ as defined in Lemma 18 in Douc et al. (2011). Neither of them requires that the $\xi_{k}$ are identically distributed. Additionally, they used their Proposition 19 to prove the claim. Note, however, that in our case we can highly simplify the proof as we have a discrete Markov chain with bounded transition probabilities. Especially, we do not even need their $\check{X}_{n}$ as in our case, their $d_{n}$ equals 1 for all $n\in\mathbb{N}.$
Lemma 5.8.
Assumption 3.1 implies that there exists a uniformly bounded function $h_{i}$ and an integer $s\in\mathbb{N}$ such that
$(i)\quad\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})$ $\xrightarrow[]{M\to\infty}1\quad\mathbb{P}_{q^{0},r^{0}}-a.s.,$
$(ii)\quad\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})$ $\xrightarrow[]{M\to\infty}0\quad\mathbb{P}_{q,r}-a.s.$
for every $(q,r)\neq(q^{0},r^{0}).$
Proof.
The proof is based on Lemma 6 in Douc et al. (2011). However, in our case $X$ is not a stationary sequence. We find an integer $s$ and bounded functions $h_{i},i=1,2,$..., such that $\mathbb{E}_{\pi}(h_{i}(X_{i},...,X_{i+s}))=1,\mathbb{E}_{\pi^{0}}(h_{i}(X_{i},...,X_{i+s}))=0$ according to the theorem from Hahn-Banach for $\pi\neq\pi^{0}$. This is a consequence of Lemma 5.1. More precisely, we choose
$h_{i}(X_{i},...,X_{i+s})=1-\frac{f_{i}(X_{i},...,X_{i+s})-1}{\mathbb{E}_{\pi^{0}}(f_{i}(X_{i},...,X_{i+s}))-1}$
with
$f_{i}(X_{i},...,X_{i+s}):=\frac{\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s})}{\mathbb{P}_{\pi}(X_{i},...,X_{i+s})}$
for $\pi\neq\pi^{0}.$ The case $\mathbb{E}_{\pi^{0}}(f_{i})=1$ cannot occur, which follows from Cauchy-Schwarz. However, the densities are uniformly bounded according to Assumption 3.1, i.e. we can also choose $h_{i}$ uniformly bounded.
Hence, we use Dolgopyat and Sarig (2023), Theorem 3.12, to prove (i).
For (ii), We define
$A_{M}:=\left\{X_{0:M}:\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})\geq\frac{1}{2}\right\}.$
It holds
$\mathbb{P}_{q,r}^{\nu}(X_{0:M}\in A_{M})$
$\quad=\sum_{k=1}^{K}\mathbb{P}_{q,r}^{\nu}(X_{0:M}\in A_{M}|Z_{0}=k)\mathbb{P}_{q,r}^{\nu}(Z_{0}=k)$
$\quad\leq\max_{k=1,...,K}\sup_{\textup{supp }\mu}\mathbb{P}^{\mu}_{q,r}\left(\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})\geq\frac{1}{2}\right)$
$\quad\leq c_{1}\exp(-M/c_{2})$
for some constant $c_{1},c_{2}>0.$ Here, we use Lemma 5.7 for the last inequality. ∎
Finally, we use Lemma 5.8 to prove Theorem 4.
Proof of Theorem 4.
The direction $\Rightarrow"$ is trivial. The other direction is similar to van Handel (2008), Theorem 7.13. First, we prove
$\ell(q,r)=\ell(q^{0},r^{0})\Rightarrow\mathbb{P}_{q^{0},r^{0}}|_{(X_{m})_{m=1,2,...}}<<\mathbb{P}_{q,r}|_{(X_{m})_{m=1,2,...}}$.
Since the number of observable states is finite, we immediately obtain this.
Now, we prove that $\mathbb{P}_{q^{0},r^{0}}\neq\mathbb{P}_{q,r}$ implies that $\mathbb{P}_{q^{0},r^{0}},\mathbb{P}_{q,r}$ are mutually singular. This is Lemma 5.8. Hence, we can then infer the claim analogously to van Handel (2008) in the proof of Theorem 7.13. ∎
5.2 Consistency
We prove Theorem 1, i.e. the consistency of the MLE. Therefore, we follow the ideas by van Handel (2008), which are stated in the homogeneous case. First, we note that both the transition probabilities and the emission probabilities are Lipschitz continuous in the parameters given Assumption 3.1. This is a direct consequence of the mean value theorem applied to $r\mapsto|e^{-rd_{m}}-e^{-r^{\prime}d_{m}}|.$
We already know that the true value is the unique maximum point of the likelihood according to Theorem 4. Therefore, we first prove
$\sup_{(q,r)\in\Theta}|\ell((X_{1},...,X_{M}),(q,r))-\ell(q,r)|\xrightarrow[]{M\to\infty}0(*).$
Based on this, we almost immediately get the claim of Theorem 1.
Proving (*) is Lemma 5.11 whose proof is divided into three steps:
  1. 1)
    Prove $\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}\left(D^{q^{0},r^{0}}_{i}\right)=\mathbb{E}\left(\ell\left(q^{0},r^{0}\right)\right).$ Since both, the transition probabilities and the emission probabilities are uniformly bounded below, we can infer (according to the majorant criterion)
    $\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}\left(D^{q^{0},r^{0}}_{i}\right)=\mathbb{E}\left(\ell\left(q^{0},r^{0}\right)\right).$
  2. 2)
    Prove $\ell((X_{1},...,X_{M}),(q,r))$ convergences for $M\to\infty$ a.s. We call the limit $\ell(q,r).$ This is Lemma 5.9
  3. 3)
    Prove the claim ()(*). This is Lemma 5.10.
Lemma 5.9 is Lemma 7.8 in van Handel (2008), which easily can be adapted to the inhomogeneous case.
Lemma 5.9.
Let Assumption 3.1 hold and let
$D_{k,\ell}^{q,r}:=\log\left(\int\mathbb{P}_{q,r}(X_{k}|Z_{k}=z)\mathbb{P}_{q,r}(Z_{k}=z|X_{k-1},...,X_{k-\ell})dz\right)$
There exist constants $\epsilon\in(0,1),\alpha$ so that
$\sup_{k\in\mathbb{N}}|D_{k,\ell}^{q,r}-D_{k}^{q,r}|\leq\alpha(1-\epsilon^{2})^{\ell}$
for all $\ell,k\in\mathbb{N}.$
Lemma 5.10 is Lemma 7.9 in van Handel (2008).
Lemma 5.10.
Let Assumption 3.1 hold. There exists a constant $\alpha$ so that
$\sup_{k\in\mathbb{N}}|D_{k}^{q,r}-D_{k}^{q^{\prime},r^{\prime}}|\leq\alpha|(q,r)-(q^{\prime},r^{\prime})|_{2}$
for all $(q,r),(q^{\prime},r^{\prime})\in\Theta.$
Lemma 5.11.
Let Assumption 3.1 hold. Then, it holds
$\sup_{q,r\in\Theta}|\ell_{M}(q,r)-\ell(q,r)|\xrightarrow[]{M\to\infty}0.$
Proof.
For simplicity, we write $\theta=(q,r)$ and $\theta^{0}=\left(q^{0},r^{0}\right).$ We start chronologically:
  1. 2)
    We apply Lemma 7.7 in van Handel (2008) to $D_{k}^{\theta}-\mathbb{E}_{\theta^{0}}(D^{\theta}_{k})$, for which we have to check whether there exists $\rho\in(0,1),C\in\mathbb{R}$ so that
    $\mathbb{E}_{\theta^{0}}\left(D_{k}^{\theta}-\mathbb{E}_{\theta^{0}}(D^{\theta}_{k})|X_{1},...,X_{\ell}\right)\leq C\rho^{k-\ell}$
    holds. Therefore, we use Lemma 5.9.
  2. 3)
    The details can be found in van Handel (2008). Let us mention that we first prove that $\ell(q,r)$ is also Lipschitz continuous. Then, the claim is a direct consequence of Lemma 5.10 and the compactness of $\Theta$.
Proof of Theorem 1.
We first show that if the likelihood has a unique maximum in $(q^{0},r^{0})$, the MLE tends to this. Therefore, we calculate $0\leq\ell(q,r)-\ell(\hat{Q}^{M},\hat{R}^{M})\xrightarrow[]{M\to\infty}0$ according to Lemma 5.11. Additionally, we know that the MLE is unique, which leads to the claim. The details are described by van Handel (2008). ∎
5.3 Central Limit Results
To prove Theorem 2, we first prove that the first derivative of $\ell((X_{1},...,X_{M}),(q^{0},r^{0}))$ is asymptotically normally distributed.
Proposition 5.12.
Let $J(q^{0},r^{0})\succ 0$ and let Assumption 3.1 hold. Then, it holds
$\lim_{M\to\infty}\frac{\nabla\ell((X_{1},...,X_{M}),(q^{0},r^{0}))}{\sqrt{M}}\xRightarrow{M\to\infty}\mathcal{N}\left(0,J^{-1}\left(q^{0},r^{0}\right)\right).$
Proof.
Let $H^{(q^{0},r^{0})}_{k}:=\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)$. First, we prove that
$\left(M^{(q^{0},r^{0})}_{n}\right)_{n\in\mathbb{N}}:=\left(\sum_{k=1}^{n}H^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(H^{(q^{0},r^{0})}_{k}|\mathcal{F}_{n-1}\right)\right)_{n\in\mathbb{N}}$
is a martingale. We just have to prove that the sum is in $L^{1}.$ However, since every summand is in $L^{1},$ we immediately obtain the claim.
Now, we check the constraints of Theorem 3.2 in Hall and Heyde (2014) to prove that $M^{(q^{0},r^{0})}_{n}/\sigma_{n}\xRightarrow{n\to\infty}\mathcal{N}(0,1)$ for
$\sigma_{n}:=\sum_{k=1}^{n}\left(\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)\right)^{2}.$
Here, the notation does not represent that $\sigma_{n}$ depends on $q^{0},r^{0}.$ Finally, we infer the claim from this asymptotic distribution. All three constraints (3.18, 3.19 and 3.20), i.e.
$\max_{k}|\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}(\nabla D^{(q^{0},r^{0})}_{k})|/\sigma_{n}\xrightarrow[]{n\to\infty}_{p}0,$
$\sum_{k=1}^{n}\left(\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)\right)^{2}/\sigma_{n}^{2}\xrightarrow[]{n\to\infty}_{p}\eta^{2}=1,$
$\sup_{n}\mathbb{E}\left(\max_{k}\frac{\left(\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)\right)^{2}}{\sigma_{n}^{2}}\right)\leq c$
for $c\in\mathbb{N},$ follow directly by the uniformly boundedness of the transition and the emission probabilities of the hidden Markov chain. By using $\sigma^{(q^{0},r^{0})}_{n}/n$ convergences to the invertible matrix $J(q^{0},r^{0})$, we can directly infer the claim.
Finally, we are ready to prove the central limit theorem.
Proof of Theorem 2.
We proceed similar to Hoadley (1971). For simplicity, we write $\theta=(q,r)$. With probability $\epsilon_{M}\xrightarrow[]{M\to\infty}1$, it holds
$0=\nabla\ell((X_{1},...,X_{M}),\hat{\theta}^{M}).$
Hence, it holds
$\nabla\ell((X_{1},...,X_{M}),\theta)|_{\theta=\hat{\theta}^{M}}-\nabla\ell((X_{1},...,X_{M}),\theta^{0})$
$=(\hat{\theta}^{M}-\theta^{0})\cdot\int_{0}^{1}\nabla^{2}\ell((X_{1},...,X_{M}),\theta^{0}+(\hat{\theta}^{M}-\theta^{0})\xi)d\xi.$
We define $I_{M}(\theta^{0}):=\int_{0}^{1}\frac{1}{M}\sum_{k=1}^{M}\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0}+(\hat{\theta}^{M}-\theta^{0})\xi)d\xi.$ Consequently,
$\sqrt{M}(\hat{\theta}^{M}-\theta^{0})I_{M}=\frac{1}{\sqrt{M}}\sum_{k=1}^{M}\nabla D_{k}((X_{1},...,X_{M}),\theta^{0}).$
We prove $I_{M}(\theta^{0})\xrightarrow[]{M\to\infty}_{p}J_{\theta^{0}}.$ It holds
$\lim_{M\to\infty}|I_{M}\left(\theta^{0}\right)-J_{\theta^{0}}|$
$=\lim_{M\to\infty}\int_{0}^{1}\frac{1}{M}\sum_{k=1}^{M}\left(\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0}+(\hat{\theta}^{M}-\theta^{0})\xi)-\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0})\right)d\xi$
$=0,$
where we used dominated convergence. Then, with Proposition 5.12 and the assumption that the matrix $J_{\theta^{0}}$ is invertible to conclude the claim. ∎
6 Discussion
From a biological perspective, the Linkage Model is useful because it explains genetic data in a relatively simple framework while accounting for linkage. From a mathematical point of view, it is particularly interesting as its stationary distribution remains the same across all markers, even though the underlying Markov chain is inhomogeneous. This specific model has already been considered in the context of Markov chains (Saloff-Coste and Zúniga, 2007). In this work, we use this property to prove limit results for the MLE in the Admixture Model.
Specifically, we investigate the consistency and central limit theorems (CLTs) of the MLE in the Linkage Model. In doing so, we prove the uniqueness of the MLE in this model. This is an important result, as in the Admixture Model—even in the supervised setting—the MLE is sometimes non-unique (Pfaffelhuber and Rohde, 2022; Heinzel et al., 2025; Heinzel, 2025).
The theoretical results can be applied in several ways, for example in marker selection, similar to Pfaff et al. (2004), who proposed using CLTs for the Admixture Model. Marker selection remains a widely studied topic (Phillips et al., 2014; Kidd et al., 2014; Xavier et al., 2022, 2020; Pfaffelhuber et al., 2020; Resutik et al., 2023; Phillips et al., 2019; Kosoy et al., 2009), and the presented results can contribute to assessing the quality of a marker set. Furthermore, the CLT represents the first published approach to quantify the uncertainty of the MLE in the Linkage Model.
Arguably, the most important application of the CLT is that they provide a theoretical foundation for a statistical test to compare the Linkage Model with the Admixture Model. This test helps determine which model better fits a given dataset. To my knowledge, this is the first data-based model selection method for the Linkage Model, as recommended by Anderson and Burnham (2004).
Of course, both the Admixture Model and the Linkage Model are simplifications of biological reality. This study only addresses the question of which of the two models fits a dataset better. It is still an open problem whether either model describes the data appropriately. Future work could include goodness-of-fit tests to evaluate whether either model is adequate at all. It would also be interesting to compare the performance of different model selection methods, such as cross-validation (Anderson and Burnham, 2004), with the statistical test developed here.
In this study, we only considered the supervised setting, i.e., the allele frequencies are assumed to be known. However, in the unsupervised setting, determining the number of ancestral populations is a major challenge in population genetics (Evanno et al., 2005; Wang, 2019; Pritchard et al., 2000; Raj et al., 2014; Verity and Nichols, 2016; Alexander and Lange, 2011). So far, none of the existing methods perform well (Garcia-Erill and Albrechtsen, 2020). A promising approach for the Linkage Model—possibly also applicable to the Admixture Model—could be methods for choosing the order of a Hidden Markov Model (HMM), such as those proposed in van Handel (2008); MacKAY (2002).
Data availability
The python scripts (application of the statistical test, evaluation of the statistical test and calculation of the variance of the MLEs) are available on GitHub (https://github.com/CarolaHeinzel/LinkageModel).
Funding
CSH is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 499552394 – SFB Small Data.
Acknowledgment
CSH thanks Peter Pfaffelhuber for his continuous and strong support.
Declaration of Conflicts of Interest
The author declares that there are no conflicts of interest.