### v2

# Highlights

Abstract
In the Admixture Model, the probability that an individual carries a certain allele at a specific marker depends on the allele frequencies in $K$ ancestral populations and the proportion of the individual’s genome originating from these populations. The markers are assumed to be independent. The Linkage Model is a Hidden Markov Model (HMM) that extends the Admixture Model by incorporating linkage between neighboring loci.
We prove consistency and asymptotic normality of maximum likelihood estimators (MLEs) for the ancestry of individuals in the Linkage Model, complementing earlier results by (Pfaff et al., 2004; Pfaffelhuber and Rohde, 2022; Heinzel, 2025) for the Admixture Model. These results are used to prove that a statistical test that allows for model selection between the Admixture Model and the Linkage Model is an asymptotic level-$\alpha$-test. Finally, we demonstrate the practical relevance of our results by applying the test to real-world data from The 1000 Genomes Project Consortium (2015).
Carola Sophia Heinzel
Department of Mathematical Stochastics,
Ernst-Zermelo Straße 1, Freiburg im Breisgau, 79140,Germany
carola.heinzel@stochastik.uni-freiburg.de
Keywords: Linkage Model, Admixture Model, Central Limit Results, Maximum Likelihood Estimator, Consistency, Statistical Test, Model Selection
1 Introduction
The Linkage Model (Falush et al., 2003) is widely used to explain the genetic data of individuals. It assumes that genetic data can be described by the ancestry proportions $q$ of an individual from $K$ ancestral populations and the allele frequencies in these populations. Furthermore, there exists a parameter $r$, which can be interpreted as the number of generations since an admixture event. The genetic distance between loci is also considered by using a Hidden Markov Model (HMM). The Admixture Model can be seen as a special case of the Linkage Model with $r=\infty$, i.e. the data is assumed to be independent across markers.
A natural question that arises is which model, the Admixture Model or the Linkage Model, fits a given dataset better. This leads to a nested model selection problem as described by e.g. Anderson and Burnham (2004). In our case, we consider the statistical hypothesis test
$H_{0}:r=\infty\quad\text{vs.}\quad H_{1}:r\in[0,\infty).$ (1)
This is a classical test problem, for which asymptotic theory exists (Wilks, 1938). However, the test is only valid as an asymptotic level-$\alpha$ test if the MLEs for both ancestry $q$ and $r$ are asymptotically normally distributed under the Linkage Model. Proving asymptotic normality requires results from the theory of HMMs.
Consistency of MLEs in HMMs with finite state and observation spaces has been addressed in foundational work by Baum and Petrie (1966) and Petrie (1969). More recent work has relaxed several of the original assumptions, e.g., (Leroux, 1992; Douc et al., 2004, 2011; Genon-Catalot and Laredo, 2006; Le Gland and Mevel, 2000a, b), but these results assume time-homogeneous Markov chains. Central limit theorems for MLEs in HMMs have also been developed (Bickel et al., 1998; Douc et al., 2004; Jensen and Petersen, 1999; Brouste and Kleptsyna, 2010). Overviews of the statistical theory of HMMs can be found in Ephraim and Merhav (2002); Cappé et al. (2005). However, the specific case of a time-inhomogeneous Markov chain that maintains the same stationary distribution for the hidden chain across all time points—as occurs in the Linkage Model when $q$ is the initial distribution—has not yet been studied.
The question whether the data should be used in STRUCTURE (Pritchard et al., 2000)(a software that estimates the ancestry and the allele frequencies for the Admixture Model or the Linkage Model) analysis has already been considered. For example, to decide whether a pair of loci is suitable for STRUCTURE, Kaeuffer et al. (2007) suggested using $r_{LD}$, a measure of linkage disequilibrium introduced by Hill and Robertson (1968). They ran STRUCTURE with the Linkage Model and used logistic regression to assess the impact of $r_{LD}$ on the detection of population structure.
To achieve our goal of establishing the theoretical properties of the test in (1), we proceed as follows: First, we precisely define both the Admixture Model and the Linkage Model. Next, we prove the asymptotic normality of the MLEs in the Linkage Model, both when the number of observations tends to infinity. Based on these results, we quantify the uncertainty of the MLEs and we construct a statistical test for (1) and prove that it is an asymptotic level-$\alpha$ test. Finally, we evaluate the statistical test through simulations and apply it to real data from The 1000 Genomes Project Consortium (2015).
2 Models
We first define the Linkage Model for one individual and bi-allelic markers in Definition 2.1. The number of alleles on chromosome $c\in\{1,...,C\}$ at marker $m\in\{1,...,M_{c}\}$ is called $X_{c,m}\in\{0,1\}.$ We denote the genetic distance in centi Morgan (cM) between the loci m1m-1 and $m$ on chromosome $c$ by $d_{c,m}$. Recall that one centi Morgan corresponds to a $0.01$ crossing over between two genetic markers per Meiosis. Let $\mathbb{S}^{K}$ be the (K1)(K-1)-dimensional simplex and let $q:=(q_{1},\ldots,q_{K})\in\mathbb{S}^{K}$ be the ancestries of the individual from population 1,,K.1,...,K. The frequency of an allele in population $k\in\{1,...,K\}$ at marker $m$, on chromosome $c$ is called $p_{c,k,m}.$ The random variable $Z_{c,m},m=1,...,M_{c},c=1,...,C$ names the ancestral population of the allele at marker $m$ on chromosome $c$. We write $(q^{0},r^{0})$ for the true values of $(q,r)$ and $\mathbb{E},\mathbb{P}$ for the expected value and the probability respectively, with respect to the true parameters $q^{0},r^{0}.$
Definition 2.1 (Linkage Model for Haploid Individuals).
We define the Markov chain
$\mathbb{P}_{q,r}(Z_{c,1}=k)$ $=q_{k},$
$\mathbb{P}_{q,r}(Z_{c,m}=\tilde{k}|Z_{c,m-1}=k)$ $=\begin{cases}e^{-d_{c,m}r}+\left(1-e^{-d_{c,m}r}\right)q_{k},\textup{ if }k=\tilde{k}\\ \left(1-e^{-d_{c,m}r}\right)q_{k},\textup{ else},\end{cases}$ (2)
$\mathbb{P}_{q,r}(Z_{c+1,1}=k|Z_{c,M_{c}})$ $=q_{k}.$
The emission probability is defined by
$\mathbb{P}_{q,r}(X_{c,m}=x|Z_{c,m})=\mathbb{P}_{q,r}\left(\textup{Ber}\left(q_{Z_{c,m}}p_{c,Z_{c,m},m}\right)=x\right).$
We define the log-likelihood, for $M_{total}:=\sum_{c=1}^{C}{M_{c}},$
$\ell((x_{1,1},...,x_{C,M}),(q,r)):=\frac{\log\left(\mathbb{P}_{q,r}((X_{c,m}=x_{c,m})_{c=1,...,C,m=1,...,M_{c}}))\right)}{M_{total}}.$
We additionally define the parameter space by $\Theta:=\{(q,r)\in[\kappa_{q},\kappa_{q}^{\prime}]^{K}\times[r_{lb},\infty]:\sum_{k=1}^{K-1}q_{k}=1\}\subset\mathbb{R}^{K+1}$.
We always assume that the allele frequencies and $(d_{c,m})_{c=1,...,C,m=1,...,M_{c}}$ are known. Definition 2.1 is for haploid individuals, which can easily be extended to the diploid case.
Remark 2.2 (Diploid Case).
Let $Z^{j}:=(Z^{j}_{c,1},...,Z^{j}_{c,M}),j=1,2,$ be two independent Markov chains with transition matrix (2). Let $X_{c,m}^{dip}$ be the number of alleles on chromosome $c,$ marker $m,$ for diploid individuals. The emission probabilities for the diploid case are defined by
$\mathbb{P}_{q,r}\left(X^{dip}_{c,m}=x|Z^{1}_{c,m},Z^{2}_{c,m}\right)$
$\hskip 28.45274pt=\begin{cases}q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m}\cdot q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m},&x=2\\ q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m}(1-q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m})+q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m}(1-q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m}),&x=1\\ (1-q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m})(1-q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m}),&x=0.\end{cases}$
For the diploid case, the information about the maternal and the paternal copies (Choi et al., 2018), i.e. phased data, is important. There are two kinds of approaches to receive phased data: laboratory-based methods (Zheng et al., 2016; Amini et al., 2014; Duitama et al., 2012) and computational methods (Choi et al., 2018). Examples for the latter type are (Snyder et al., 2015; Loh et al., 2016; Delaneau et al., 2012). Falush et al. (2003) also proposed a method to deal with linkage without phasing. However, in this study, we assume that we have phased data as e.g. in The 1000 Genomes Project Consortium (2015).
While phased data is important for the Linkage Model, the Admixture Model can deal with unphased data without any problems. It is defined in Definition 2.3.
Definition 2.3 (Admixture Model).
The Admixture Model is a special case of the Linkage Model with $r=\infty.$ Especially, this means that it holds
$X_{c,m}\sim\textup{Bin}(2,\langle q_{\cdot}^{0},p_{c,\cdot,m}\rangle).$
Extending the models to a general number of individuals and markers with arbitrary number of alleles is straightforward.
Finally, we define the statistical test that corresponds to test problem 2.
Definition 2.4 (Statistical Test).
We define the test statistic for the test (1) by
$\Lambda:=-2\ln\left(\frac{\max\left\{\ell((X_{1,1},...,X_{C,M}),(q,\infty)):q\in\mathbb{S}^{K}\right\}}{\max\{\ell((X_{1,1},...,X_{C,M}),(q,r)):(q,r)\in\Theta\}}\right).$
Furthermore, let $\chi^{2}_{1-\alpha}$ be the $1-\alpha$-quantile of the $\chi^{2}(1)$-distribution. We reject $H_{0}$, if $\Lambda>\chi^{2}_{1-\alpha}.$
3 Theoretical Main Results
In this chapter, we prove the main results, i.e. consistency and central limit results, if the number of markers, $M_{total},$ tends to infinity. Therefore, we assume that Assumption 3.1 holds.
Assumption 3.1.
We assume that $C=1$ and that we have haploid individuals. We additionally assume that
  • (A1)
    it holds $1>\kappa_{p}^{\prime}\geq p_{\cdot,m}\geq\kappa_{p}>0$ for all $m\in\{1,...,M\}.$
  • (A2)
    it holds $p_{k,m}\neq p_{\ell,m}$ for infinitely many markers m and for all $k,\ell\in\{1,..,K\},k\neq\ell.$
  • (A3)
    Let $\lambda_{c,m}$ be the second largest eigen value of the transition matrix $(\mathbb{P}(Z_{c,m}=k|Z_{c,m-1}=k^{\prime}))_{k=1,...,K,k^{\prime}=1,...,K}$. It holds
    $\prod_{c=1}^{C}\prod_{m=1}^{M_{c}}\lambda_{c,m}\xrightarrow[]{M_{total}\to\infty}0.$
  • (A4)
    there exists a constant $\kappa_{d}>0$, such that $d_{c,m}\geq\kappa_{d}$ for all $m\in\{1,...,M_{c}\},c=1,...,C.$
  • (A5)
    It holds $d_{m}\leq d_{up}$ for all $m\in\{1,...,M\}.$
By the definition of $\Theta$, we already assumed that $q\in[\kappa_{q},\kappa^{\prime}_{q}]$ and $r\geq r_{lb}.$
Figure 1 gives an overview of the results in this paper.
Unique MLEAssumption 3.1ConsistencyAssumption 3.1CLTAssumption 3.1,invertible Fischer Information Asymptotic Distributionof the Test StatisticAssumption 3.1
Figure 1: Overview of results in this section and their requirements.
We also state the theory under the following conditions concerning the data.
Remark 3.2 (Assumptions).
We assume that the data is haploid and that we only have bi-allelic markers. Additionally, we only consider one individual. However, extending the theory to more general cases is straightforward.
We start with the consistency of the MLE (Theorem 1).
Theorem 1 (Consistency of the MLE).
Let Assumption 3.1 hold. Then, for the MLE
$\left(\hat{Q}^{C,M},\hat{R}^{C,M}\right):=\textup{argmax}\{(q,r)\mapsto\ell((x_{1,1},...,x_{C,M}),(q,r))\}$
it holds
$\mathbb{P}\left(\lim_{M_{total}\to\infty}\bigg|\left(\hat{Q}^{C,M},\hat{R}^{C,M}\right)-\left(q^{0},r^{0}\right)\bigg|\geq\epsilon\right)=0$
for any $\epsilon>0.$
Based on Theorem 1, we can infer a CLT for the MLE, if the true parameters $\left(q^{0},r^{0}\right)$ are in the interior of the parameter space.
Theorem 2 (Central Limit Theorem for the MLE).
Let Assumption 3.1 holdand let the Fischer information
$J(q^{0},r^{0}):=\lim_{M_{total}\to\infty}\mathbb{E}\left(\frac{\partial^{2}}{\partial(q,r)^{2}}\ell\left((x_{1,1},...,x_{C,M}),(q,r))\big|_{(q,r)=\left(q^{0},r^{0}\right)}\right)\right)$
be positive definite. Let the true parameters be in the interior of the parameter space. Then, it holds
$\sqrt{M_{total}}\left((\hat{Q}^{C,M},\hat{R}^{C,M})-(q^{0},r^{0})\right)\xRightarrow{M_{total}\to\infty}\mathcal{N}\left(0,J(q^{0},r^{0})^{-1}\right).$
Remark 3.3.
Douc (2005) has proven the invertibility of the Fischer Information under certain constraints for stationary HMMs, i.e. in our case, if the allele frequencies and the distance between the markers are identical for every marker. This also means that the markers are either all on different chromosomes or on the same chromosome.
According to Wilks (1938), since the MLE in the Linkage Model is asymptotically normally distributed, Theorem 3 about the asymptotic distribution of the test statistic holds.
Theorem 3 (Asymptotic distribution of the test statistic).
It holds
$\Lambda\xRightarrow{M_{total}\to\infty}\chi^{2}(1)$
under the null hypothesis of (1).
4 Application to Data
In this section, we first evaluate the performance of the test from Definition 2.4 by using simulated data. Afterwards, we apply the test to data from The 1000 Genomes Project Consortium (2015) and compare the uncertainty of the MLEs in the Linkage Model to the ones in the Admixture Model.
4.1 Simulated Data
To evaluate the statistical test in terms of Type 1 and Type 2 errors, I simulated data under both the Linkage Model and the Admixture Model for different values of $M$, $r$, and the marker distances $(d_{c,m})_{c=1,\dots,C,\ m=1,\dots,M_{c}-1}$. Specifically, I simulated 100 markers on a single chromosome with $d_{1,m}=d_{1,m+1}=:d$ for all $m\in\{1,...,M-1\}$, where $d\in\{0.1,0.5,1,2,5,10\}$ and $r\in\{1,10,100\}$. The significance level was set to $0.05$.
I computed the Type 1 and Type 2 error rates by repeating the experiment 100 times for each possible combination of $r$ and $d$. The results are shown in Figure 2. Notably, the Type 1 error remains small and consistently below the $0.05$ threshold. Moreover, both the Type 1 and Type 2 errors tend to increase with increasing values of $r$ and $d$. This behavior is expected, as larger values of $r$ and $d$ imply greater similarity between the two models, making them harder to distinguish.
In summary, the simulation results indicate that the test exhibits high power while maintaining a low Type 1 error rate.
Figure 2: Evaluation of the statistical test by using simulated data for different values of $r$ and $d$.
4.2 Real Data
I also applied the statistical test to the data from The 1000 Genomes Project Consortium (2015) with the ancestry informative marker (AIM) set by Kidd et al. (2014) consisting of $M_{total}=55$ markers on 20 different autosomes ($C=20$). For the analysis, individuals from five populations were included: Africa (AFR), Europe (EUR), South Asia (SAS), East Asia (EAS), and Admixed Americans (AMR), i.e., $K=5$. We always used all 2504 individual except the one that we are testing to calculate the allele frequencies as relative frequency. The values for $d$ for the AIM set by Kidd et al. (2014) in cM are shown in Figure 3.
Figure 3: Genetic Distances of the markers in the AIM set by Kidd et al. (2014) rounded to integer values.
The distance between loci on a single chromosome ranged from below 1 cM to approximately 129 cM.
The results of the statistical test are shown in Figure 4. In total, for 312 out of 2504 individuals (approximately 12.5%), the null hypothesis — that the Admixture Model provides a better fit to the data — cannot be rejected. Notably, there are substantial differences across populations: for individuals from Africa and East Asia, the proportion for which the null hypothesis is rejected is considerably lower compared to other groups. This highlights that it is not sufficient to define a general rule for choosing between the Linkage Model and the Admixture Model based solely on the marker set. Instead, a method — such as the statistical test defined in equation (1) — is needed, which allows model selection on an individual or population-specific basis.
Figure 4: Results of the statistical test from Definition 2.4 for the data from The 1000 Genomes Project Consortium (2015).
We further compare the covariance matrices of the MLEs obtained under the Linkage Model and the Admixture Model for individual HG00096 from Great Britain. For this analysis, I again used the AIM set introduced by Kidd et al. (2014). The results are shown in Figure 6 and Figure 6, corresponding to the Admixture and Linkage Models, respectively.
As expected, the variance of the estimators is higher in the Linkage Model due to the larger number of parameters. While both models clearly identify EUR as the primary ancestral population, the estimated admixture proportions differ significantly between models (estimated ancestry of almost 1 from Europe in the Admixture Model and estimated ancestry of approximately 0.8 from Europe in the Linkage Model).
Figure 5: Covariance Matrix for the MLE in the Admixture Model. We considered individual HG00096. The MLE for $q$ was $(1,0,0,0,0).$
Figure 6: Covariance Matrix of the MLE in the Linkage Model. We considered individual HG00096. The MLE for $q$ was $(0.8,0.04,0.04,0.08,0.04)$ and for $r$ approximately $1.8.$
Remark 4.1 (Test for a whole population).
We also provide the code for testing whether the Linkage Model or the Admixture Model fits better to the data of $N$ individuals on GitHub. In this framework, two modeling approaches are possible: either (i) the recombination rate $r$ is assumed to vary across individuals, requiring the estimation of $N\times K$ parameters, or (ii) a single global recombination rate $r$ is assumed, reducing the number of parameters to $N\times(K-1)+1$. We chose the latter approach (ii) to reduce model complexity.
5 Proofs of the Main Results
We first prove that the MLE $\left(\hat{Q}^{M},\hat{R}^{M}\right):=\left(\hat{Q}^{1,M},\hat{R}^{1,M}\right)$ is asymptotically unique. Based on this, we prove consistency and central limit results for the MLE in the Linkage Model. For all proofs, we assume $C=1$ to simplify the notation.
To prove the consistency and the CLT, we represent the log-likelihood as a sum, i.e. it holds
$\ell((X_{1},...,X_{M}),(q,r))$
$\quad=\frac{1}{M}\sum_{m=1}^{M}\underbrace{\log\left(\int\mathbb{P}_{q,r}(X_{m}|Z_{m}=z)\mathbb{P}_{q,r}(Z_{m}=dz|X_{1},...,X_{m-1})\right)}_{=:D_{m}^{q,r}}$
according to van Handel (2008), Proposition 6.4, for the homogeneous case. This can easily extended to the inhomogeneous case.
5.1 Unique MLE
Proving the uniqueness of a MLE for HMM has been already considered extensively for a finite state space, see e.g. (Finesso, 1990; Gilbert, 1959; Blackwell and Koopmans, 1957; Petrie, 1969). In this section, we aim to prove the asymptotic uniqueness of the MLE in the Linkage Model, i.e. we aim to prove Theorem 4. Therefore, we use that the limit
$\ell((X_{1},...,X_{M}),(q,r))\xrightarrow[]{M\to\infty}\ell(q,r)$ (3)
exists $\mathbb{P}$-almost surely. We will prove this in section 5.2.
Theorem 4 (Unique Maximum Likelihood Estimator).
Let Assumption 3.1 hold. Then, it holds
$\ell(q^{0},r^{0})>\ell(q,r)\Leftrightarrow q,r\neq q^{0},r^{0}.$ (4)
To prove Theorem 4, we first prove that the parameters are identifiable based on a finite number of observations (Lemma 5.1). Based on this, we prove that if $\ell(q^{0},r^{0})\neq\ell(q,r)$, then the measures $\mathbb{P}_{q,r}|_{(X_{m})_{m=1,2,...}}$ and $\mathbb{P}_{q^{0},r^{0}}|_{(X_{m})_{m=1,2,...}}$ are mutually different. Then, we can conclude the claim.
We start by proving the identifiability of the parameter for $K=2.$
Lemma 5.1 (Identifiability of the Parameters).
Let $K=2$ and let $d_{1},d_{2}>0.$ If the allele frequencies of $p_{\cdot,i}$ and $p_{\cdot,i+2}$ are linearly independent and in addition $p_{1,m}\neq p_{2,m}\forall m\in\{i,i+1,i+2\}$, the parameters $(q,r)$ are identifiable based on the random variables $X_{i},...,X_{i+2}.$
Proof.
To save some notation, we write $A_{i}(z_{i},z_{i+1})=\mathbb{P}(Z_{i+1}=z_{i+1}|Z_{i}=z_{i}),B_{j}(x_{j},z_{j})=\mathbb{P}(X_{j}=x_{j}|Z_{j}=z_{j}).$ In order to apply Kruskal’s Theorem, we first need a decomposition of the marginal density of $X_{1},X_{2},X_{3}.$ Let $M_{j}$ be a stochastic matrix, where the $i^{th}$ row is denoted by $m^{j}_{i}.$ We define
$[M_{1},M_{2},M_{3}]_{x_{1},x_{2},x_{3}}=\sum_{i=1}^{2}m^{1}_{i}(x_{1})m^{2}_{i}(x_{2})m^{3}_{i}(x_{3})$
and $M_{1},M_{2},M_{3}]$ is a three-dimensional array for which the element $(x_{1},x_{2},x_{3})$ equals $M_{1},M_{2},M_{3}]_{x_{1},x_{2},x_{3}}.$ Without loss of generality, we assume $i=1$. It holds
$\mathbb{P}(X_{1}$ $=x_{1},X_{2}=x_{2},X_{3}=x_{3})$
$=\sum_{z_{1},z_{2},z_{3}}\pi_{z_{1}}A_{1}(z_{1},z_{2})A_{2}(z_{2},z_{3})B_{2}(x_{2},z_{2})B_{3}(x_{3},z_{3})B_{1}(x_{1},z_{1})$
$=[\pi B_{1}A_{1},B_{2},\underbrace{A_{2}B_{3}}_{=:M_{2}}]_{x_{1},x_{2},x_{3}}.$
The proof is divided into two steps:
  • (i)
    prove that the entries of $M_{1}:=B_{1}A_{1},M_{2},B_{2}$ are identifiable from the marginal density of $X_{1},X_{2},X_{3}$. Therefore, we use theory by Allman et al. (2009); Kruskal (1977).
  • (ii)
    conclude that the parameters $q,r$ are identifiable.
We start with (i). Therefore, we have to prove $rk(M_{1})=rk(B_{2})=rk(M_{2})=2.$ It holds
$\det(B_{2})=\bigg|\begin{pmatrix}q_{1}p_{1,2}&(1-q_{1})p_{2,2}\\ 1-q_{1}p_{1,2}&1-(1-q_{1})p_{2,2}\end{pmatrix}\bigg|=q_{1}p_{1,2}-(1-q_{1})p_{2,2}.$
Hence, $B_{2}$ has the full rang if and only if $q_{1}\neq\frac{p_{1,2}}{p_{1,2}+p_{2,2}}.$ We have
$M_{1}=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$
with entries
$a$ $=\mathbb{P}(X_{1}=0|Z_{2}=1)$
$=\sum_{z_{1}=1}^{2}\mathbb{P}(X_{0}=0|Z_{1}=z_{1})\mathbb{P}(Z_{1}=z_{1}|Z_{2}=1)$
$=(1-q_{1}p_{1,1})\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right)+\left(1-(1-q_{1})p_{2,1}\right)(1-e^{-rd_{1}})(1-q_{1})$
$=(1-q_{1}p_{1,1})(q_{1}+e^{-rd_{1}}(1-q_{1}))+(1-p_{2,1}+q_{1}p_{2,1})(1-q_{1})(1-e^{-rd_{1}})$
$b$ $=\mathbb{P}(X_{1}=1|Z_{2}=1)$
$=q_{1}p_{1,1}\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right)+p_{2,1}(1-q_{1})^{2}(1-e^{-rd_{1}}),$
$c$ $=\mathbb{P}(X_{1}=0|Z_{2}=2)$
$=(1-q_{1}p_{1,1})(1-e^{-rd_{1}})(1-q_{1})+\left(1-(1-q_{1})p_{2,1}\right)\left((1-e^{-rd_{1}})(1-q_{1})+e^{-rd_{1}}\right),$
$d$ $=\mathbb{P}(X_{1}=1|Z_{2}=2)$
$=q_{1}p_{1,1}(1-e^{-rd_{1}})(1-q_{1})+p_{2,1}(1-q_{1})\left((1-e^{-rd_{1}})(1-q_{1})+e^{-rd_{1}}\right).$
It holds
$\det(M_{1})=(2e^{-2rd_{1}}q_{1}^{2}-3e^{-2rd_{1}}q_{1}+e^{-2rd_{1}}-4e^{-rd_{1}}q_{1}^{2}+6e^{-rd_{1}}q_{1}-3e^{-rd_{1}}+2q_{1}^{2}-3q_{1}+1)$
$\quad\quad\quad\quad\quad\quad\cdot(p_{1,1}q_{1}+p_{2,1}q_{1}-p_{2,1})=0$
$\Leftrightarrow q_{1}=\frac{p_{2,1}}{p_{1,1}+p_{2,1}}$
$\quad\quad\lor(2e^{-2rd_{1}}q_{1}^{2}-3e^{-2rd_{1}}q_{1}+e^{-2rd_{1}}-4e^{-rd_{1}}q_{1}^{2}+6e^{-rd_{1}}q_{1}-3e^{-rd_{1}}+2q_{1}^{2}-3q_{1}+1)=0.$
It holds
$2e^{-2rd_{1}}q_{1}^{2}-3e^{-2rd_{1}}q_{1}+e^{-2rd_{1}}-4e^{-rd_{1}}q_{1}^{2}+6e^{-rd_{1}}q_{1}-3e^{-rd_{1}}+2q_{1}^{2}-3q_{1}+1=0$
$\Leftrightarrow q_{1}=\frac{e^{-rd_{1}}}{e^{-rd_{1}}-1}\lor q_{1}=\frac{e^{-rd_{1}}-3}{2(e^{-rd_{1}}-1)}.$
Fortunately, it holds $\frac{e^{-rd_{1}}}{e^{-rd_{1}}-1}\leq 0$ and $\frac{e^{-rd_{1}}-3}{2(e^{-rd_{1}}-1)}\geq 3/2.$ Hence, we do not have to take care of these terms.
We additionally calculate
$\mathbb{P}(X_{k-2}$ $=x\mid Z_{k-1}=z)$
$=\sum_{z^{\prime}=1}^{2}\mathbb{P}(X_{k-2}=x,Z_{k-2}=z^{\prime}\mid Z_{k-1}=z)$
$\propto\sum_{z^{\prime}=1}^{2}\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=z^{\prime})\cdot\mathbb{P}(Z_{k-2}=z^{\prime}\mid Z_{k-1}=z)$
$\propto\underbrace{\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=0)}_{=(q_{1}p_{1,3})^{x}(1-q_{1}p_{1,3})^{1-x}}\cdot\mathbb{P}(Z_{k-1}=0\mid Z_{k-2}=z)$
$\hskip 28.45274pt+\underbrace{\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=1)}_{=(q_{2}p_{2,3})^{x}(1-q_{2}p_{2,3})^{1-x}}\cdot\underbrace{\mathbb{P}(Z_{k-1}=1\mid Z_{k-2}=z)}_{=1_{z=1}e^{-rd_{2}}+q_{z}(1-e^{-rd})}$
$\propto(q_{1}p_{1,3})^{x}(1-q_{1}p_{1,3})^{1-x}\left(1_{z=1}e^{-rd_{2}}+q_{z}(1-e^{-rd})\right)$
$\hskip 28.45274pt+(q_{2}p_{2,3})^{x}(1-q_{2}p_{2,3})^{1-x}\left(1_{z=2}e^{-rd_{2}}+q_{z}(1-e^{-rd})\right).$
This leads to
$M_{2}=\begin{pmatrix}a_{2}&b_{2}\\ c_{2}&d_{2}\end{pmatrix}$
with entries
$a_{2}$ $=(1-q_{1}p_{1,3})\left(e^{-rd_{2}}+q_{1}(1-e^{-rd_{2}})\right)+(1-(1-q_{1})p_{2,3})q_{1}(1-e^{-rd_{2}})$
$b_{2}$ $=(q_{1}p_{1,3})\left(e^{-rd_{2}}+q_{1}(1-e^{-rd_{2}})\right)+(1-q_{1})p_{2,3}q_{1}(1-e^{-rd_{2}}),$
$c_{2}$ $=(1-q_{1}p_{1,3})(1-q_{1})(1-e^{-rd_{2}})$
$\quad\quad+(1-(1-q_{1})p_{2,3})\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right),$
$d_{2}$ $=q_{1}p_{1,3}(1-q_{1})(1-e^{-rd_{2}})+(1-q_{1})p_{2,3}\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right).$
It holds
$\det(M_{2})=e^{-rd_{2}}(p_{1,3}q_{1}+p_{2,3}q_{1}-p_{2,3})=0$
$\Leftrightarrow r=\infty\lor q_{1}=\frac{p_{2,3}}{p_{1,3}+p_{2,3}}.$
For $r=\infty,$ it holds $\mathbb{P}(X_{1}=x)=(q_{1}p_{1,1}+(1-q_{1})p_{2,1})^{x}(1-(q_{1}p_{1,1}+(1-q_{1})p_{2,1}))^{1-x}.$ Then, $q$ is identifiable if and only if $p_{2,1}\neq p_{1,1}.$
Now, we come to step (ii), i.e. we have to ensure that the parameters are uniquely determined based on the entries. For the emission probabilities, this is a direct consequence of the assumption $p_{k,m}\neq p_{\ell,m}.$ Then, since $d_{m}>0,$ it is also immediate to infer the value $r$ from the other parameters.
In the latter, we will – without a proof – just assume that the parameters are identifiable from the marginal density of the observations. The proof is basically the same. However, we of course have to increase the number of observations form $3$ for $K=2$ to 2(K1)+2.2(K-1)+2. Additionally, we then have to verify again that the matrices of the decomposition of the marginal density have the full rank under certain conditions.
Now, we aim to infer (4) from the identifiability. Therefore, we first prove a law of large numbers. Then, based on this, we conclude (4).
To prove Theorem 4, we will use the following particularity of the hidden chain.
Lemma 5.2 (Invariant Measure).
For each $m\in\mathbb{N},$ the probability measure $q$ is the unique invariant measure for the transition matrix of $Z_{m}$.
Additionally, we will use that the hidden chain is uniformly ergodic (Lemma 5.4), which is defined as follows.
Definition 5.3 (Uniformly Ergodic).
Let $|\cdot|_{TV}$ be the total variation norm. A Markov chain with transition matrices $T_{i}^{\theta}$ is called uniformly ergodic, if it holds
$|T^{\theta}_{n}...T^{\theta}_{1}-\pi|_{TV}\leq c_{\theta}\rho_{\theta}^{n}$
for $\rho_{\theta}<1,c_{\theta}\in\mathbb{R}.$
Lemma 5.4 (Uniformly Ergodicity in the Linkage Model).
Under Assumption 3.1, $(Z_{m})_{m=1,2,...}$ is uniformly ergodic.
Proof.
According to Saloff-Coste and Zúniga (2007), Theorem 3.3, the convergence result holds from Lemma 5.2. ∎
Lemma 5.5.
Assumption 3.1 implies that there exists a function $h_{i}$ and an integer $s\in\mathbb{N}$ such that
$(i)\quad\frac{1}{n}\sum_{i=1}^{n}h_{i}(X_{i},...,X_{i+s})$ $\xrightarrow[]{n\to\infty}1\quad\mathbb{P}_{(q^{0},r^{0})}\,-a.s.$
$(ii)\quad\frac{1}{n}\sum_{i=1}^{n}h_{i}(X_{i},...,X_{i+s})$ $\xrightarrow[]{n\to\infty}0\quad\mathbb{P}_{(q,r)}\,-a.s.$
for every $(q,r)\neq(q^{0},r^{0}).$
Proof.
The proof is based on Lemma 6 in Douc et al. (2011). The only difference is that in our case $X$ is not a stationary sequence. We find an integer $s$ and a bounded functions $h_{i},i=1,2,$..., such that $\mathbb{E}_{\pi}(h_{i}(X_{i},...,X_{i+s}))=1,\mathbb{E}_{\pi^{0}}(h_{i}(X_{i},...,X_{i+s}))=0$ according to the theorem from Hahn-Banach for $\pi\neq\pi^{0}$. More precisely, we choose $h_{i}=1-\frac{f_{i}-1}{\mathbb{E}_{\pi^{0}}(f_{i})-1}$ with $f_{i}:=\frac{\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s})}{\mathbb{P}_{\pi}(X_{i},...,X_{i+s})}$ for $\theta\neq\theta^{0}$. The case $\mathbb{E}_{\pi^{0}}(f_{i})=1$ cannot occur according to Lemma 5.1. However, we have to ensure that there exists a constant $c_{q,r,q^{0},r^{0}}>0,$ such that $\mathbb{E}_{\pi^{0}}(f_{i})\geq 1+c_{q,r,q^{0},r^{0}}.$ It holds
$\mathbb{E}_{\pi^{0}}(f_{i})$ $=1+\int\frac{(\mathbb{P}_{\pi}(X_{i},...,X_{i+s})-\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s}))^{2}}{\mathbb{P}_{\pi}(X_{i},...,X_{i+s})}dX_{i}....X_{i+s}$
$\geq 1+\frac{1}{\kappa^{2(s+1)}}\int(\mathbb{P}_{\pi}(X_{i},...,X_{i+s})-\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s}))^{2}\,dX_{i}....X_{i+s}$
$\geq 1+\frac{1}{\kappa^{2(s+1)}K^{s+1}}|\mathbb{P}_{\pi}(X_{i},....,X_{i+s})-\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s})|_{TV}$
for $\kappa=\kappa_{q}\lor\kappa_{p}.$ There exists a $\delta>0$, such that $|\mathbb{P}_{\pi}(X_{m}=x|Z_{m}=z)-\mathbb{P}_{\pi^{0}}(X_{m}=x|Z_{m}=z)|\geq\delta$ for all $z_{m},z,x,z_{m-1}$, the total variation distance between $\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s})$ and $\mathbb{P}_{\pi}(X_{i},...,X_{i+s})$ is bounded below.
The densities are uniformly bounded according to Assumption 3.1, i.e. $h_{i}$ is uniformly bounded.
Hence, we proceed similar to the proof of Theorem 3.12 in Dolgopyat and Sarig (2023) to prove (i) and (ii). Dolgopyat and Sarig (2023) considered the convergence of $\sum_{i=1}^{n}f_{i}(Z_{i},Z_{i+1})$ under certain conditions. Here, we consider the convergence of $\sum_{i=1}^{n}f_{i}(g_{i}(Z_{i}),g_{i+1}(Z_{i+1}))$ under certain conditions. The only difference compared to the proof in Dolgopyat and Sarig (2023) is that we have to prove
$\|\mathbb{E}(f_{i}(g_{i}(Z_{i}),g_{i+1}(Z_{i+1}))-\mathbb{E}(f_{i}(g_{i}(Z_{i}),g_{i+1}(Z_{i+1}))|Z_{j})\|_{2}\leq c\rho^{i-j}$
for $i>j.$ However, this claim is a direct consequence of Lemma 5.4.
We use Lemma 5.5 to prove the identifiably of the parameters. This is stated in the following theorem (Theorem 7.13 in van Handel (2008)).
Proof of Theorem 4.
The direction $\Rightarrow"$ is trivial. The other direction is similar to van Handel (2008), Theorem 7.13. First, we prove
$\ell(q,r)=\ell(q^{0},r^{0})\Rightarrow\mathbb{P}_{q^{0},r^{0}}|_{(X_{m})_{m=1,2,...}}<<\mathbb{P}_{q,r}|_{(X_{m})_{m=1,2,...}}$.
Therefore, we use Shiryaev (2016). Since the number of observable states is finite, the claim is a direct consequence of this.
Now, we prove that $\mathbb{P}_{q^{0},r^{0}}\neq\mathbb{P}_{q,r}$ implies that $\mathbb{P}_{q^{0},r^{0}},\mathbb{P}_{q,r}$ are mutually singular. This is Lemma 5.5. ∎
5.2 Consistency
We prove Theorem 1, i.e. the consistency of the MLE. Therefore, we follow the ideas by van Handel (2008), which are stated in the homogeneous case. First, we note that both the transition probabilities and the emission probabilities are Lipschitz continuous in the parameters.
Lemma 5.6.
Let $d_{m}\leq d_{up}$ for all $m\in\{1,...,M\}.$ There exists two constants $L_{e}\in\mathbb{R}$ and $L_{t}\in\mathbb{R}$ such that
$|\mathbb{P}_{q,r}(X_{m}=x|Z_{m}=z_{m})-\mathbb{P}_{q,r}(X_{m}=x|Z_{m}=z_{m})|$ $\leq L_{e}|(q,r)-(q^{\prime},r^{\prime})|$
$|\mathbb{P}_{q,r}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})-\mathbb{P}_{q^{\prime},r^{\prime}}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})|$ $\leq L_{t}|(q,r)-(q^{\prime},r^{\prime})|.$
Proof.
It holds
$|\mathbb{P}_{q,r}(X_{m}=x|Z_{m}=z_{m})-\mathbb{P}_{q^{\prime},r^{\prime}}(X_{m}=x|Z_{m}=z_{m})|$ $=|q^{\prime}_{z_{m}}-q_{z_{m}}|p_{z_{m},m}.$
Hence, we can choose $L_{e}=1.$ It also holds
$|\mathbb{P}_{q,r}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})-\mathbb{P}_{q^{\prime},r^{\prime}}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})|$
$=|(1-e^{-rd_{m}})q_{k}-(1-e^{-r^{\prime}d_{m}})q^{\prime}_{k}|$
$=\big|(1-e^{-rd_{m}})(q_{k}-q^{\prime}_{k})+q^{\prime}_{k}(e^{-r^{\prime}d_{m}}-e^{-rd_{m}})\big|$
$\leq|1-e^{-rd_{m}}|\;|q_{k}-q^{\prime}_{k}|+|q^{\prime}_{k}|\,|e^{-rd_{m}}-e^{-r^{\prime}d_{m}}|$
$\leq|q_{k}-q^{\prime}_{k}|+|e^{-rd_{m}}-e^{-r^{\prime}d_{m}}|.$
Moreover, by the mean value theorem there exists $\xi\in[r,r^{\prime}]$ such that
$|e^{-rd_{m}}-e^{-r^{\prime}d_{m}}|=d_{m}\,|r-r^{\prime}|\,e^{-\xi d_{m}}\leq d_{m}\,|r-r^{\prime}|.$
We already know that the true value is the unique maximum point of the likelihood according to Theorem 4. Hence, we first prove
$\sup_{(q,r)\in\Theta}|\ell((X_{1},...,X_{M}),(q,r))-\ell(q,r)|\xrightarrow[]{M\to\infty}0$ (5)
under Assumption 3.1. Based on this, we almost immediately get the claim of Theorem 1.
Proving (5) is Lemma 5.9 whose proof is divided into three steps:
  1. 1)
    Prove $\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}\left(D^{q^{0},r^{0}}_{i}\right)=\mathbb{E}\left(\ell\left(q^{0},r^{0}\right)\right).$
  2. 2)
    Prove $\ell((X_{1},...,X_{M}),(q,r))$ converges for $M\to\infty$ a.s. We call the limit $\ell(q,r).$ This is Lemma 5.7
  3. 3)
    Prove equation (5). This is Lemma 5.9.
Lemma 5.7 and Lemma 5.8 are Lemma 7.8 and Lemma 7.9 in van Handel (2008), respectively, and can easily be adapted to the inhomogeneous case.
Lemma 5.7.
Let Assumption 3.1 hold and let
$D_{k,\ell}^{q,r}:=\log\left(\int\mathbb{P}_{q,r}(X_{k}|Z_{k}=z)\mathbb{P}_{q,r}(Z_{k}=z|X_{k-1},...,X_{k-\ell})dz\right)$
There exists a constant $\alpha$ so that
$\sup_{k\in\mathbb{N}}|D_{k,\ell}^{q,r}-D_{k}^{q,r}|\leq\alpha(1-\epsilon^{2})^{\ell}$
for all $\ell,k\in\mathbb{N}.$
Lemma 5.8.
Let Assumption 3.1 hold. There exists a constant $\alpha$ so that
$\sup_{k\in\mathbb{N}}|D_{k}^{q,r}-D_{k}^{q^{\prime},r^{\prime}}|\leq\alpha|(q,r)-(q^{\prime},r^{\prime})|_{2}$
for all $(q,r),(q^{\prime},r^{\prime})\in\Theta.$
Lemma 5.9 is stated for the homogeneous case as Proposition 7.5 in van Handel (2008).
Lemma 5.9.
Let Assumption 3.1 hold. Then, equation (5) holds.
Proof.
For simplicity, we write $\theta=(q,r)$ and $\theta^{0}=\left(q^{0},r^{0}\right).$ We start chronologically:
  1. 1)
    Since both, the transition probabilities and the emission probabilities are uniformly bounded below, we infer (according to the majorant criterion)
    $\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}\left(D^{q^{0},r^{0}}_{i}\right)=\mathbb{E}\left(\ell\left(q^{0},r^{0}\right)\right).$
  2. 2)
    We apply Lemma 7.7 in van Handel (2008) to $D_{k}^{\theta}-\mathbb{E}_{\theta^{0}}(D^{\theta}_{k})$, for which we have to check whether there exists $\rho\in(0,1),C\in\mathbb{R}$ so that
    $\mathbb{E}_{\theta^{0}}\left(D_{k}^{\theta}-\mathbb{E}_{\theta^{0}}(D^{\theta}_{k})|X_{1},...,X_{\ell}\right)\leq C\rho^{k-\ell}$
    holds. Therefore, we use Lemma 5.7.
  3. 3)
    The details can be found in van Handel (2008). Let us mention that we first prove that $\ell(q,r)$ is also Lipschitz continuous by using Lemma 5.6. Then, the claim is a direct consequence of Lemma 5.8 and the compactness of $\Theta$.
Proof of Theorem 1.
We first show that if the likelihood has a unique maximum in $(q^{0},r^{0})$, the MLE tends to this. Therefore, we calculate $0\leq\ell(q,r)-\ell(\hat{Q}^{M},\hat{R}^{M})\xrightarrow[]{M\to\infty}0$ according to Lemma 5.9. Additionally, we know that the MLE is unique, which leads to the claim. The details are described by van Handel (2008), Theorem 7.6. ∎
5.3 Central Limit Results
To prove the asymptotic normality of the MLE (Theorem 2), we first prove that the first derivative of $\ell((X_{1},...,X_{M}),(q^{0},r^{0}))$ is asymptotically normally distributed. Based on this, we conclude the claim with similar techniques as in Hoadley (1971).
Proposition 5.10.
Let $J(q^{0},r^{0})\succ 0$ and let Assumption 3.1 hold. Then, it holds
$\lim_{M\to\infty}\frac{\nabla\ell((X_{1},...,X_{M}),(q^{0},r^{0}))}{\sqrt{M}}\xRightarrow{M\to\infty}\mathcal{N}\left(0,J^{-1}_{q^{0},r^{0}}\right).$
Proof.
The proof idea is the same as the one of Proposition 5 in (Douc and Matias, 2001). We first define
$N_{k}^{(q^{0},r^{0})}:=\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)$.
Let $\mathcal{F}_{n}$ be the sigma field that is generated by $\nabla D^{(q^{0},r^{0})}_{k},k=1,...,n$. We notice that
$\left(M^{(q^{0},r^{0})}_{n}\right)_{n\in\mathbb{N}}:=\left(\sum_{k=1}^{n}N_{k}^{(q^{0},r^{0})}+\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{n}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{n}\right)|\mathcal{F}_{n-1}\right)\right)_{n\in\mathbb{N}}$
is a martingale.
Now, we check the constraints of Theorem 3.2 in Hall and Heyde (2014) to prove that $M^{(q^{0},r^{0})}_{n}/\sigma_{n}\xRightarrow{n\to\infty}\mathcal{N}(0,1)$ for
$\sigma_{n}:=\sum_{k=1}^{n}\left(N_{k}^{(q^{0},r^{0})}\right)^{2}.$
Here, the notation omits that $\sigma_{n}$ depends on $q^{0},r^{0}.$
All three constraints (i.e. equations 3.18, 3.19 and 3.20 in Hall and Heyde (2014)), i.e.
$\max_{k}|\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}(\nabla D^{(q^{0},r^{0})}_{k})|/\sigma_{n}\xrightarrow[]{n\to\infty}_{p}0,$
$\sum_{k=1}^{n}\left(\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)\right)^{2}/\sigma_{n}^{2}\xrightarrow[]{n\to\infty}_{p}\eta^{2}=1,$
$\sup_{n}\mathbb{E}\left(\max_{k}\frac{\left(\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)\right)^{2}}{\sigma_{n}^{2}}\right)\leq c$
for $c\in\mathbb{R},$ follow directly by the uniformly boundedness of the transition and the emission probabilities of the hidden Markov chain. By using $\sigma^{(q^{0},r^{0})}_{n}/n$ convergences to the invertible matrix $J(q^{0},r^{0})$, we can directly infer the claim.
Finally, we are ready to prove the asymptotic normality of $\hat{Q}^{M},\hat{R}^{M}$.
Proof of Theorem 2.
We proceed similar to Hoadley (1971). For simplicity, we write $\theta=(q,r)$. With probability $\epsilon_{M}\xrightarrow[]{M\to\infty}1$, it holds
$0=\nabla\ell((X_{1},...,X_{M}),\hat{\theta}^{M}).$
Hence, we get
$\nabla\ell((X_{1},...,X_{M}),\theta)|_{\theta=\hat{\theta}^{M}}-\nabla\ell((X_{1},...,X_{M}),\theta^{0})$
$=(\hat{\theta}^{M}-\theta^{0})\cdot\int_{0}^{1}\nabla^{2}\ell((X_{1},...,X_{M}),\theta^{0}+(\hat{\theta}^{M}-\theta^{0})\xi)d\xi.$
We define $I_{M}(\theta^{0}):=\int_{0}^{1}\frac{1}{M}\sum_{k=1}^{M}\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0}+(\hat{\theta}^{M}-\theta^{0})\xi)d\xi.$ Consequently,
$\sqrt{M}(\hat{\theta}^{M}-\theta^{0})I_{M}(\theta^{0})=\frac{1}{\sqrt{M}}\sum_{k=1}^{M}\nabla D_{k}((X_{1},...,X_{M}),\theta^{0}).$
We prove $I_{M}(\theta^{0})\xrightarrow[]{M\to\infty}_{p}J\left(\theta^{0}\right).$ It holds
$\lim_{M\to\infty}|I_{M}\left(\theta^{0}\right)-J\left(\theta^{0}\right)|$
$=\lim_{M\to\infty}\int_{0}^{1}\frac{1}{M}\sum_{k=1}^{M}\left(\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0}+(\hat{\theta}^{M}-\theta^{0})\xi)-\mathbb{E}\left(\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0})\right)\right)d\xi$
$=0,$
where we used dominated convergence. Then, with Proposition 5.10 and the assumption that the matrix $J\left(\theta^{0}\right)$ is invertible, we conclude the claim. ∎
6 Discussion
From a biological perspective, the Linkage Model is useful because it explains genetic data in a relatively simple framework while accounting for linkage. From a mathematical point of view, it is particularly interesting as the stationary distribution of the hidden Markov chain remains the same across all markers, even though the underlying Markov chain is inhomogeneous. This specific model has already been considered in the context of Markov chains (Saloff-Coste and Zúniga, 2007). In this work, we use this property to prove limit results for the MLE in the Linkage Model.
Specifically, we investigate the consistency and the asymptotic normality of the MLE in the Linkage Model. In doing so, we prove the uniqueness of the MLE in this model. This is an important result, as in the Admixture Model—even in the supervised setting—the MLE is sometimes non-unique (Pfaffelhuber and Rohde, 2022; Heinzel et al., 2025; Heinzel, 2025).
The theoretical results can be applied in several ways, for example in marker selection, similar to Pfaff et al. (2004), who proposed using CLTs for the Admixture Model. Marker selection remains a widely studied topic (Phillips et al., 2014; Kidd et al., 2014; Xavier et al., 2022, 2020; Pfaffelhuber et al., 2020; Resutik et al., 2023; Phillips et al., 2019; Kosoy et al., 2009), and the presented results can contribute to assessing the quality of a marker set. Furthermore, the CLT represents the first published approach to quantify the uncertainty of the MLE in the Linkage Model.
Arguably, the most important application of the CLT is that it provides a theoretical foundation for a statistical test to compare the Linkage Model with the Admixture Model. This test helps determine which model better fits a given dataset. To my knowledge, this is the first data-based model selection method for the Linkage Model.
Of course, both the Admixture Model and the Linkage Model are simplifications of biological reality. This study only addresses the question of which of the two models fits a dataset better. It is still an open problem whether either model describes the data appropriately. Future work could include goodness-of-fit tests to evaluate whether either model is adequate at all. It would also be interesting to compare the performance of different model selection methods, such as cross-validation (Anderson and Burnham, 2004), with the statistical test developed here.
In this study, we only considered the supervised setting, i.e., the allele frequencies are assumed to be known. However, in the unsupervised setting, determining the number of ancestral populations is a major challenge in population genetics (Evanno et al., 2005; Wang, 2019; Pritchard et al., 2000; Raj et al., 2014; Verity and Nichols, 2016; Alexander and Lange, 2011). So far, none of the existing methods perform well (Garcia-Erill and Albrechtsen, 2020). A promising approach for the Linkage Model—possibly also applicable to the Admixture Model—could be methods for choosing the order of a HMM, such as those proposed in van Handel (2008); Mackay (2002).
Data availability
The python scripts (application of the statistical test, evaluation of the statistical test and calculation of the variance of the MLEs) are available on GitHub (https://github.com/CarolaHeinzel/LinkageModel).
Funding
CSH is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 499552394 – SFB Small Data.
Acknowledgment
CSH thanks Peter Pfaffelhuber for his continuous and strong support.
Declaration of Conflicts of Interest
The author declares that there are no conflicts of interest.